# A Pollak Proof for the Number of Weakly Increasing Parking Functions

### 02 Apr 2026 v2

@@ Abstract.
1. Introduction
@@
Parking functions were first studied by Pyke (1959) and by Konheim and Weiss (1966); they independently showed that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$ for all $n\geq 1$ (OEIS, , A000169). Riordan (1969) credits Pollak with a different and particularly elegant proof of this result based on a circular-street argument; refer to (Martínez Mori, 2024) for an expository description of Pollak’s technique.
Parking functions are also closely related to Catalan combinatorics; refer to Stanley (2015) for a comprehensive survey of this area. In particular, a parking function $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ is said to be weakly increasing if $a_{1}\leq a_{2}\leq\cdots\leq a_{n}$. Let $\mathrm{PF}_{n}^{\uparrow}\subseteq\mathrm{PF}_{n}$ denote the set of weakly increasing parking functions of length $n\geq 1$. Then, it is well-known that $|\mathrm{PF}_{n}^{\uparrow}|=C_{n}=\frac{1}{n+1}\binom{2n}{n}$, where $C_{n}$ is the $n$th Catalan number (OEIS, , A000108). This result can be established in various ways, most notably directly through the recursive relation $C_{n}=\sum_{i=1}^{n}C_{i}C_{n-i}$ for $n\geq 1$ with $C_{0}=1$ or bijectively through another known Catalan object (most typically Dyck paths, as in (Armstrong et al., 2016, Section 2.2)).
@@
2. A Circular-Street Argument for Weakly Increasing Parking Functions
@@
This follows as a direct application of the balls and bars method. Specifically, consider the placement of $n$ bars and $n$ balls on a straight line from left to right. In this order, the $i$th ball corresponds to the $i$th car, whose preference is $1$ plus the total number of bars to its left. Therefore, the choice of $n$ places for the bars from the total of $2n$ different places fully determines the cars’ weakly increasing preferences. ∎
@@
Figure 1. Illustrations accompanying Example 1.
@@
Consider $4$ cars parking on a circular one-way street with $5$ spots, i.e., after spot $5$, the street loops back into spot $1$. The weakly increasing preferences of length $4$ are in bijection with the contents of length $5$ via the content map $\kappa$ whose inverse is the weakly increasing word map $\tau$. For example, $\kappa\left((1,1,3,4)\right)=(2,0,1,1,0)$ and $\tau\left((2,0,1,1,0)\right)=(1,1,3,4)$.
The circular shift operation on tuples, denoted by $\omega$, partitions the set of contents into equivalence classes of size $5$. In particular, the elements of the orbit
@@
also of size $5$, and with exactly one (not necessarily weakly increasing) parking function: $(3,4,1,1)$. Through $\kappa$, these preference rotations are in bijection with the equivalence class of contents. For example,
@@
Finally, the spot ultimately left unoccupied is invariant under $\phi$, as noted by the dashed arrows and the unlabeled parking outcomes. For example, the spot ultimately left unoccupied under both $(3,4,1,1)$ and $(1,1,3,4)$ is spot $5$. Since $(3,4,1,1)$ is the unique parking function in its equivalence class of preference rotations, $(1,1,3,4)$ is the unique parking function in its equivalence class of weakly increasing preferences (shaded in golden in Figure 1).
3. Discussion

### 09 Apr 2026 v3

@@ Abstract
keywords:
cyclic shifts, Catalan numbers, parking functions
1 Introduction
@@
Parking functions were first studied by Pyke (1959) and by Konheim and Weiss (1966); they independently showed that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$ for all $n\geq 1$ (OEIS Foundation Inc., , A000169). Riordan (1969) credits Pollak with a different and particularly elegant proof of this result based on a circular-street argument; refer to (Martínez Mori, 2024) for an expository description of Pollak’s technique.
Parking functions are also closely related to Catalan combinatorics; refer to Stanley (2015) for a comprehensive survey of this area. In particular, a parking function $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ is said to be weakly increasing if $a_{1}\leq a_{2}\leq\cdots\leq a_{n}$. Let $\mathrm{PF}_{n}^{\uparrow}\subseteq\mathrm{PF}_{n}$ denote the set of weakly increasing parking functions of length $n\geq 1$. Then, it is well-known that $|\mathrm{PF}_{n}^{\uparrow}|=C_{n}=\frac{1}{n+1}\binom{2n}{n}$, where $C_{n}$ is the $n$th Catalan number (OEIS Foundation Inc., , A000108). This result can be established in various ways, most notably directly through the recursive relation $C_{n}=\sum_{i=1}^{n}C_{i}C_{n-i}$ for $n\geq 1$ with $C_{0}=1$ or bijectively through another known Catalan object (most typically Dyck paths, as in (Armstrong et al., 2016, Section 2.2)).
@@
2 A Circular-Street Argument for Weakly Increasing Parking Functions
@@
This follows as a direct application of the balls and bars method. Specifically, consider the placement of $n$ bars and $n$ balls on a straight line from left to right. In this order, the $i$th ball corresponds to the $i$th car, whose preference is 11 plus the total number of bars to its left. Therefore, the choice of $n$ places for the bars from the total of $2n$ different places fully determines the cars’ weakly increasing preferences. ∎
@@
Figure 1: Illustrations accompanying Example 4.
@@
Consider 44 cars parking on a circular one-way street with 55 spots, i.e., after spot 55, the street loops back into spot 11. The weakly increasing preferences of length 44 are in bijection with the contents of length 55 via the content map $\kappa$ whose inverse is the weakly increasing word map $\tau$. For example, $\kappa\left((1,1,3,4)\right)=(2,0,1,1,0)$ and $\tau\left((2,0,1,1,0)\right)=(1,1,3,4)$.
The circular shift operation on tuples, denoted by $\omega$, partitions the set of contents into equivalence classes of size 55. In particular, the elements of the orbit
@@
also of size 55, and with exactly one (not necessarily weakly increasing) parking function: $(3,4,1,1)$. Through $\kappa$, these preference rotations are in bijection with the equivalence class of contents. For example,
@@
Finally, the spot ultimately left unoccupied is invariant under $\phi$, as noted by the dashed arrows and the unlabeled parking outcomes. For example, the spot ultimately left unoccupied under both $(3,4,1,1)$ and $(1,1,3,4)$ is spot 55. Since $(3,4,1,1)$ is the unique parking function in its equivalence class of preference rotations, $(1,1,3,4)$ is the unique parking function in its equivalence class of weakly increasing preferences (shaded in golden in Figure 1).
3 Discussion